OpenStudy (anonymous):

Gain a medal! What are the zeros of the polynomial function: f(x)= x^3− x^2 − 12x? 0, –4, –3 0, –4, 3 0, 4, –3 0, 4, 3

OpenStudy (lucaz):

first factor out x

OpenStudy (lucaz):

you get x(x^2-x-12) you can again factor the quadratic equation

OpenStudy (lucaz):

sorry, the quadratic expression

OpenStudy (lucaz):

@r4ge you know how to factor quadratics?

OpenStudy (anonymous):

no I don't :( @lucaz

OpenStudy (lucaz):

the quadratic is x^2-x-12 we need two values that add up to -1 (from the -x term) and that when multiplied we get -12 (from the -12 term)

OpenStudy (lucaz):

let's try -4 and 3, -4+3=-1, -4*3=-12 they work, so: x^2-x-12 is the same as (x-4)(x+3)

OpenStudy (lucaz):

so, the original expression becomes: x(x-4)(x+3), the roots are the values of x that make the function f(x)=0 if we plug in x=0 we will get 0*(0-4)(0+3), this is equal to zero

OpenStudy (lucaz):

f(4) = 4*(4-4)*(4+3) = 4*0*(7) = also equal to zero

OpenStudy (lucaz):

f(-3) = -3*(-3-4)*(-3+3) = -3*(-7)*0 = zero

OpenStudy (lucaz):

the values that make f(x) = 0 are 0, 4 and -3

OpenStudy (anonymous):

thank you sososo much!

OpenStudy (lucaz):

no problem